OpenStudy (anonymous):

Find the next three terms of 9, 3, 1,...

hartnn (hartnn):

notice the pattern. every next term is 1/3 times the previous, could u observe now ?

OpenStudy (anonymous):

umm so it goes up 1/3?

hartnn (hartnn):

next term =1/3 is correct. what about other 2 ? whats 1/3 (1/3) =... ?

OpenStudy (anonymous):

1/9?

hartnn (hartnn):

yes! next term = 1/3 (1/9) =... ?

OpenStudy (anonymous):

ok i got it thanks :)

hartnn (hartnn):

welcome ^_^